# Exploiting Diffusion Prior for Generalizable Pixel-Level Semantic Prediction

### v1

@@ Abstract
Contents generated by recent advanced Text-to-Image (T2I) diffusion models are sometimes too imaginative for existing off-the-shelf property semantic predictors to estimate due to the immitigable domain gap. We introduce DMP, a pipeline utilizing pre-trained T2I models as a prior for pixel-level semantic prediction tasks. To address the misalignment between deterministic prediction tasks and stochastic T2I models, we reformulate the diffusion process through a sequence of interpolations, establishing a deterministic mapping between input RGB images and output prediction distributions. To preserve generalizability, we use low-rank adaptation to fine-tune pre-trained models. Extensive experiments across five tasks, including 3D property estimation, semantic segmentation, and intrinsic image decomposition, showcase the efficacy of the proposed method. Despite limited-domain training data, the approach yields faithful estimations for arbitrary images, surpassing existing state-of-the-art algorithms. The code is available at https://github.com/shinying/dmp.
Figure 1: Generalized pixel-level semantic prediction. (left) We leverage the pre-trained text-to-image diffusion model  [46] as a prior for various pixel-level semantic prediction tasks. (right) With only a small amount of labeled training data in a limited domain (i.e., 10K bedroom images with labels) for each task, our method performs favorably against SOTA predictors [25, 5, 15] on arbitrary images.
@@ 1 Introduction
Text-to-image (T2I) diffusion models [46, 49, 18, 11] have achieved unprecedented progress on text-guided image generation, producing highly imaginative and realistic images from diverse and free-from textual descriptions. These advancements open up a new era of AI-aided content creation with applications spanning various domains [21, 42, 52, 60, 1, 48, 44]. However, the creativity of images generated by T2I models poses challenges for off-the-shelf property (e.g., depth, normal, segmentation) prediction methods [5, 25, 15] due to the domain gap. For example, the ZoeDepth [5] approach fails to accurately predict the depth of the cubism painting shown in the sixth column in Figure 1. Such property predictions are vital for understanding high-level semantics of generated contents and can facilitate various downstream applications such as 3D imaging [53] and relighting [70]. We term these property prediction tasks as pixel-level semantic prediction in this paper.
@@
Existing pixel-level semantic prediction models are typically trained on “real-world” images regardless of the training dataset scale. While these models aim for generalization, bridging the domain gap between real-world and T2I-generated images remains challenging, as we demonstrate in the right-hand side of Figure 1. Several recent efforts have been made to take advantage of the inherent generalizability of pre-trained T2I models to solve various image-to-image translation tasks [42, 21, 7]. Although the pixel-level semantic prediction tasks can also be formulated as image-to-image translation, we show in Section 4 that these methods are not directly applicable to the pixel-level semantic prediction problems.
@@
Leveraging pre-trained T2I models as a prior for the pixel-level semantic prediction tasks is challenging for two reasons. First, most pixel-level semantic prediction tasks are inherently deterministic, posing difficulties when adapting a pre-trained T2I model designed for stochastic text-to-image generation. Second, striking a balance between learning target tasks and retaining the inherent generalizability of pre-trained T2I models is crucial. In other words, learned pixel-level semantic predictors should generalize to arbitrary images from the training data in a limited domain.
@@
In this paper, we propose DMP (Diffusion Models as Priors) to leverage the pre-trained T2I model [46] as a prior for generalized pixel-level semantic prediction. To resolve the determinism-stochasticity misalignment, we introduce a deterministic mapping between the input RGB images and output prediction distributions. Specifically, we reformulate the diffusion process as a chain of interpolations between input RGB images and their corresponding output signals, where the importance of input images gradually increases over the diffusion process. The reverse diffusion (i.e., known as denoising or generation in original T2I) process becomes a series of transformations that progressively synthesize desired output signals from input images. Without randomization, such as Gaussian noise imposed, the mapping is entirely deterministic. In addition, to retain the generalizability of the pre-trained T2I model while learning target tasks, we use low-rank adaptation [23] to fine-tune the pre-trained model with the aforementioned deterministic diffusion process for each pixel-level semantic prediction task. Figure 1 demonstrates the generalization ability of the proposed method on the deterministic normal, depth, and segmentation prediction problems.
@@
We conduct extensive quantitative and qualitative experiments on five pixel-level semantic prediction tasks to evaluate the proposed DMP approach: 3D property estimation (depth, normal), semantic segmentation, and intrinsic image decomposition (albedo, shading). We show that with only a small amount of limited-domain training data (i.e., 10K bedroom training images with labels), the proposed method can provide faithful estimations of the in-domain and unseen images, especially those that the existing SOTA algorithms struggle to handle effectively. We summarize the contributions as follows:
@@
  • We propose DMP, an approach leveraging the pre-trained T2I model as a prior for pixel-level semantic prediction tasks.
@@
  • We design a diffusion process that adapts the stochastic T2I model for deterministic pixel-level semantic prediction problems.
@@
  • We use five pixel-level semantic prediction tasks to validate that the proposed method obtains faithful estimation on arbitrary images despite training with a small amount of data in a limited domain.
@@ 2 Related Work
Diffusion models. Diffusion models estimate a target data distribution by modeling the transition from a noisy version of the distribution [56, 22]. Recently, diffusion models have shown unprecedented quality in the text-to-image [46, 49] setting by training on large-scale datasets [51]. The advancements unleash various text-guided image manipulation applications [21, 42, 52, 1, 44, 9, 7, 39]. The stochasticity in the generation process, a preferred property in most applications, derives from the initial noise and additional noise added during the denoising process. However, the pixel-level semantic prediction problems are usually deterministic. Some recent methods propose to adopt deterministic sampling algorithms [57, 29, 34] and reformulate the diffusion and generative process by $\alpha$-blending and de-blending [20]. Though the mapping between initial noise and outputs in target distributions is deterministic, the correlation between each noise and output sample is stochastic. If adapted to deterministic tasks, the model may generate high-quality outputs unfaithful to input images.
@@
Image-to-image translation. The task aims to learn a mapping between two visual domains. Early efforts [75, 24, 30, 76, 41, 8] mostly make use of generative adversarial networks and cycle consistency loss to learn the translation from scratch. With StyleGAN [27, 28] showing high-quality synthesis on certain categories, some methods [62, 45] propose to achieve image-to-image translation using pre-trained StyleGAN by training an additional encoder. However, the translation is limited to certain categories. Recently, following the success of large-scale text-to-image models, attempts have been made to perform image-to-image translation with pre-trained diffusion models [21, 42, 7, 64]. These methods, however, are not directly applicable to the pixel-level semantic prediction task.
@@
Fine-tuning text-to-image diffusion models. In addition to image-to-image translation, pre-trained diffusion models are adopted to take additional modalities [72, 32], be customized on certain objects [16, 47] or styles [73], synthesize videos [35, 19, 17], etc. These methods train additional zero-initialized layers [72], manipulate attention modules [17, 32], learn a token embedding [16, 73], or learn parameter offsets with low-rank matrices [23]. In this work, we adopt the low-rank adaptation [23] to fine-tune only parameter offsets of attention layers.
@@
Generative prior for semantic prediction. Prior work has leveraged pre-trained generative models as a prior for other tasks, such as representation learning [13], as the latent features of pre-trained generative models are found to be rich in semantics [63, 69, 67]. Bhattad et al. [6] manipulates style latents of StyleGAN [26] and reveals its learned ability to estimate image properties, but the generalizability is limited. Other works reuse [2, 68] or merge [61] latent features of denoising models in pre-trained text-to-image diffusion models to perform segmentation. In this work, instead of developing a specific approach for a particular task, we focus on analyzing the potential of pre-trained models as a prior for general semantic prediction through a universal transferring framework.
@@ 3 Method
Our goal is to leverage the pre-trained T2I diffusion model as a prior to learn a pixel-level semantic prediction task from a set of labeled training data $\mathcal{D}=\{(x^{i},y^{i})\}^{n}_{i=1}$, where $x^{i}$ denotes the input image, $y^{i}$ indicates the corresponding output (e.g., depth map), and $x^{i},y^{i}\in\mathbb{R}^{H\times W\times 3}$. We first describe the text-to-diffusion models used as the prior in Section 3.1. We then introduce the proposed DMP approach in Section 3.2.
@@ 3.1 Text-to-Image Diffusion Models
We use the pre-trained T2I latent diffusion model [46] as the prior. It consists of an autoencoder and a U-Net. The autoencoder converts between an image $y\in\mathbb{R}^{H\times W\times 3}$ and its latent feature $\tilde{y}\in\mathbb{R}^{h\times\ w\times c}$, where $(h,w)=(H/8,W/8)$ and $c$ represents the channel size of latent features. Since we do not modify the autoencoder in the proposed approach, we use $y$ to represent the latent feature $\tilde{y}$ for simplifying the annotation in this paper.
@@ 3.2 Leveraging Diffusion Prior
Figure 2: Deterministic diffusion process. We formulate the diffusion process as a chain of interpolations between an input image $x$ and output $y$. The U-Net model is fine-tuned to gradually transform the input $x$ to the desired pixel-level semantic prediction $y$.
Figure 3: 3D property estimation of arbitrary input images. The first row shows the input image, while the remaining rows present the normal and depth estimated by different approaches. The proposed DMP method gives faithful estimation, even on the images where the off-the-shelf [25, 5] schemes fail to handle.
Deterministic diffusion. The diffusion process described in Eq. (1) is designed specifically for stochastic image generation. However, the mapping between input images and outputs in pixel-level semantic prediction problems is typically deterministic. We observe that directly applying the diffusion process in Eq. (1) to the semantic prediction tasks introduces unnecessary variation in outputs that leads to apparent artifacts. Therefore, we use the blending strategy [20, 33] and re-design the diffusion process as follows. Instead of converting from noise maps to images in the conventional T2I method, the diffusion process in our DMP approach directly maps between the input image $x$ and output $y$. As illustrated in Figure 2, the proposed diffusion process is formulated as
@@
Parameterization. We explore various parameterizations for the U-Net model to make different predictions. As discussed in Section 4.5, we empirically find that the v-prediction [50] works well for the pixel-level semantic prediction tasks. Specifically, the U-Net model $v_{\theta}$ is fine-tuned using the following mean square loss:
@@ 4.1 Experiment Setup
We evaluate the proposed DMP approach using five pixel-level prediction tasks, including 3D property estimation (i.e., depth and normal), semantic segmentation, and intrinsic image decomposition (i.e., albedo and shading).
@@
In-domain Out-of-domain
Normal Depth Normal Depth
@@
SPADE [40] 0.0708 0.1635 0.2132 0.4961 0.1379 0.1268 0.2833 0.3587 0.3190 0.2554
DRIT++ [31] 0.0784 0.1723 0.3792 0.2458 0.2134 0.1350 0.3006 0.4373 0.2585 0.3216
SDEdit [37] 0.2599 0.5087 0.4656 0.3533 0.3240 0.2675 0.5293 0.6640 0.2495 0.3382
DDIB [59] 0.1849 0.4210 0.3087 0.5130 0.2367 0.2271 0.4847 0.6275 0.2788 0.3120
IP2P (hard) [7] 0.3017 0.5468 0.4834 0.3235 0.3358 0.3168 0.5757 0.6450 0.2252 0.3461
IP2P (learned) [7] 0.3550 0.7181 0.3965 0.3302 0.3494 0.3397 0.6836 0.5182 0.2664 0.3261
VISII [39] 0.2081 0.4386 0.3498 0.4405 0.2912 0.2448 0.4895 0.5364 0.2855 0.3181
DMP 0.0514 0.1156 0.1072 0.8861 0.1020 0.0872 0.1886 0.2117 0.6395 0.1360
@@
Datasets. We first generate diverse text descriptions using a large language model [38] by filling a keyword in a prompt template modified from the one used in Parmar et al. [43]. We then use a text-to-image diffusion model [46] to synthesize the images according to the text descriptions. Second, we use the following off-the-shelf predictors to generate the pseudo ground truth for each image: Omnidata v2 [25] for surface normal, ZoeDepth [5] for monocular depth, EVA-2 [15] for semantic segmentation, and PIE-Net [12] for intrinsic image decomposition (albedo, shading). Finally, we follow the protocol used in Bhattad et al. [6] to generate a set of training data, and three sets of test data:
@@
  • VISII [39]: We use their approach for finetuning with the training set and inference.
@@
In-domain Out-of-domain
Albedo Shading Albedo Shading
SPADE [40] 0.0021 0.0031 0.0030 0.0040
DRIT++ [31] 0.0296 0.0309 0.0392 0.0408
SDEdit [37] 0.0375 0.0501 0.0471 0.0671
DDIB [59] 0.0411 0.0403 0.0443 0.0557
@@
IP2P (learned) [7] 0.0215 0.0290 0.0250 0.0309
VISII [39] 0.0145 0.0275 0.0246 0.0285
DMP 0.0041 0.0051 0.0064 0.0070
Table 4: Effect of different parameterizations. We fine-tune the U-Net model to predict different signals in each reverse diffusion step to get the final output: predicting input image $x$, predicting output $y$, and v-prediction described in Eq. (3). The experiment is conducted on the surface normal prediction task.
@@
Figure 6: Qualitative comparisons between various parameterizations. We fine-tune the U-Net model to predict different signals in each reverse diffusion step to get the final output: predicting input image $x$, predicting output $y$, and v-prediction described in Eq. (3).

### v2

# Exploiting Diffusion Prior for Generalizable Dense Prediction

@@ Abstract
Contents generated by recent advanced Text-to-Image (T2I) diffusion models are sometimes too imaginative for existing off-the-shelf dense predictors to estimate due to the immitigable domain gap. We introduce DMP, a pipeline utilizing pre-trained T2I models as a prior for dense prediction tasks. To address the misalignment between deterministic prediction tasks and stochastic T2I models, we reformulate the diffusion process through a sequence of interpolations, establishing a deterministic mapping between input RGB images and output prediction distributions. To preserve generalizability, we use low-rank adaptation to fine-tune pre-trained models. Extensive experiments across five tasks, including 3D property estimation, semantic segmentation, and intrinsic image decomposition, showcase the efficacy of the proposed method. Despite limited-domain training data, the approach yields faithful estimations for arbitrary images, surpassing existing state-of-the-art algorithms. The code is available at https://github.com/shinying/dmp.
Figure 1: Generalized dense prediction. (left) We leverage the pre-trained text-to-image diffusion model  [57] as a prior for various dense prediction tasks. (right) With only a small amount of labeled training data in a limited domain (i.e., 10K bedroom images with labels) for each task, our method performs favorably against SOTA predictors [29, 7, 17] on arbitrary images.
@@ 1 Introduction
Text-to-image (T2I) diffusion models [57, 61, 21, 13] have achieved unprecedented progress on text-guided image generation, producing highly imaginative and realistic images from diverse and free-from textual descriptions. These advancements open up a new era of AI-aided content creation with applications spanning various domains [24, 49, 64, 72, 2, 59, 52]. However, the creativity of images generated by T2I models poses challenges for off-the-shelf dense (e.g., depth, normal, segmentation) prediction methods [7, 29, 17] due to the domain gap. For example, the ZoeDepth [7] approach fails to accurately predict the depth of the cubism painting shown in the sixth column in Figure 1. Such property predictions are vital for understanding high-level semantics of generated contents and can facilitate various downstream applications such as 3D imaging [65] and relighting [84].
@@
Existing dense prediction models are typically trained on “real-world” images regardless of the training dataset scale. While these models aim for generalization, bridging the domain gap between real-world and T2I-generated images remains challenging, as we demonstrate in the right-hand side of Figure 1. A potential solution is to take advantage of the inherent generalizability of pre-trained T2I models, for example, by formulating dense prediction as image-to-image translation. However, while several recent efforts have been made to solve various image-to-image translations with pre-trained T2I models [49, 24, 9], we show in Section 4 that these methods are not directly applicable.
@@
Leveraging pre-trained T2I models as a prior for dense prediction is challenging for two reasons. First, most dense prediction tasks are inherently deterministic, posing difficulties when adapting a pre-trained T2I model designed for stochastic text-to-image generation. Second, it is crucial to strike a balance between learning target tasks and retaining the inherent generalizability of pre-trained T2I models. In other words, learned dense predictors should generalize to arbitrary images from the training data in a limited domain.
@@
In this paper, we propose DMP (Diffusion Models as Priors) to leverage the pre-trained T2I model [57] as a prior for generalized dense prediction. To resolve the determinism-stochasticity misalignment, we introduce a deterministic mapping between the input RGB images and output prediction distributions. Specifically, we reformulate the diffusion process as a chain of interpolations between input RGB images and their corresponding output signals, where the importance of input images gradually increases over the diffusion process. The reverse diffusion (i.e., known as denoising or generation in original T2I) process becomes a series of transformations that progressively synthesize desired output signals from input images. Without randomization, such as Gaussian noise imposed, the mapping is entirely deterministic. In addition, to retain the generalizability of the pre-trained T2I model while learning target tasks, we use low-rank adaptation [26] to fine-tune the pre-trained model with the aforementioned deterministic diffusion process for each dense prediction task. Figure 1 demonstrates the generalization ability of the proposed method on the deterministic normal, depth, and segmentation prediction problems.
@@
We conduct extensive quantitative and qualitative experiments on five dense prediction tasks to evaluate the proposed DMP approach: 3D property estimation (depth, normal), semantic segmentation, and intrinsic image decomposition (albedo, shading). We show that with only a small amount of limited-domain training data (i.e., 10K bedroom training images with labels), the proposed method can provide faithful estimations of the in-domain and unseen images, especially those that the existing SOTA algorithms struggle to handle effectively. We summarize the contributions as follows:
@@
  • We propose DMP, an approach leveraging the pre-trained T2I model as a prior for dense prediction tasks.
@@
  • We design an image-to-prediction diffusion process that adapts the stochastic T2I model for deterministic dense prediction problems.
@@
  • We use five dense prediction tasks to validate that the proposed method obtains faithful estimation on arbitrary images despite training with a small amount of data in a limited domain.
@@ 2 Related Work
Diffusion models. Diffusion models estimate a target data distribution by modeling the transition from a noisy version of the distribution [68, 25]. Recently, diffusion models have shown unprecedented quality in the text-to-image [57, 61] setting by training on large-scale datasets [63]. The advancements unleash various text-guided image manipulation applications [24, 49, 64, 2, 52, 11, 9, 46]. The stochasticity in the generation process, a preferred property in most applications, derives from the initial noise and additional noise added during the denoising process. However, the dense prediction problems are usually deterministic. Some recent methods adopt deterministic sampling algorithms [69, 33, 41] and reformulate the diffusion and generative process by $\alpha$-blending and de-blending [23]. Though the mapping between initial noise and outputs in target distributions is deterministic, the correlation between each noise and output sample is stochastic. If adapted to deterministic tasks, the model may generate high-quality outputs unfaithful to input images.
@@
Image-to-image translation. The task aims to learn a mapping between two visual domains. Early efforts [90, 28, 35, 91, 48, 10] mostly make use of generative adversarial networks and cycle consistency loss to learn the translation from scratch. With StyleGAN [31, 32] showing high-quality synthesis on certain categories, some methods [74, 55] seek to achieve image-to-image translation using pre-trained StyleGAN by training an additional encoder. However, the translation is limited to certain categories. Recently, following the success of large-scale text-to-image models, attempts have been made to perform image-to-image translation with pre-trained diffusion models [24, 49, 9, 76]. These methods, however, are not directly applicable to the dense prediction task.
@@
Fine-tuning text-to-image diffusion models. In addition to image-to-image translation, pre-trained diffusion models are adopted to take additional modalities [86, 39], be customized on certain objects [18, 58] or styles [87], and synthesize videos [42, 22, 19]. These methods train additional zero-initialized layers [86], manipulate attention modules [19, 39], learn a token embedding [18, 87], or learn parameter offsets with low-rank matrices [26]. In this work, we adopt the low-rank adaptation [26] to fine-tune only parameter offsets of attention layers.
@@
Generative prior for dense prediction. Prior work has leveraged pre-trained generative models as a prior for other tasks, such as representation learning [15], as the latent features of pre-trained generative models are found to be rich in semantics [75, 83, 81]. Bhattad et al. [8] manipulates style latents of StyleGAN [30] and reveals its learned ability to estimate image properties, but the generalizability is limited. Some works reuse [3, 82, 88, 20] or merge [73] latent features of pre-trained denoising U-Nets to perform segmentation and depth estimation. Others [79, 78] transform generation models into multi-task generalists by standardizing the outputs of tasks as images. In this work, instead of developing a specific approach for a particular task, we focus on analyzing the potential of pre-trained models as a prior for general dense prediction through a universal transferring framework.
@@ 3 Method
Our goal is to leverage the pre-trained T2I diffusion model as a prior to learn a dense prediction task from a set of labeled training data $\mathcal{D}=\{(x^{i},y^{i})\}^{n}_{i=1}$, where $x^{i}$ denotes the input image, $y^{i}$ indicates the corresponding output (e.g., depth map), and $x^{i},y^{i}\in\mathbb{R}^{H\times W\times 3}$. We first describe the text-to-diffusion models used as the prior in Section 3.1. Then we introduce the proposed DMP approach in Section 3.2.
@@ 3.1 Text-to-Image Diffusion Models
We use the pre-trained T2I latent diffusion model [57] as the prior. It consists of an autoencoder and a U-Net. The autoencoder converts between an image $y\in\mathbb{R}^{H\times W\times 3}$ and its latent feature $\tilde{y}\in\mathbb{R}^{h\times\ w\times c}$, where $(h,w)=(H/8,W/8)$ and $c$ represents the channel size of latent features. Since we do not modify the autoencoder in the proposed approach, we use $y$ to represent the latent feature $\tilde{y}$ to simplify the annotation in this paper.
@@ 3.2 Leveraging Diffusion Prior
Figure 2: Deterministic diffusion process. We formulate the diffusion process as a chain of interpolations between an input image $x$ and output $y$. The U-Net model is fine-tuned to gradually transform the input $x$ to the desired dense prediction $y$.
Figure 3: 3D property estimation of arbitrary input images. The first row shows the input images, while the remaining rows present the normals and depth estimated by different approaches. The proposed DMP method gives faithful estimation, even on the images where the off-the-shelf [29, 7] schemes fail to handle.
Deterministic diffusion. The diffusion process described in Eq. (1) is designed specifically for stochastic image generation. However, the mapping between input images and outputs in dense prediction problems is typically deterministic. We observe that directly applying the diffusion process in Eq. (1) to the dense prediction tasks introduces unnecessary variation in outputs that leads to apparent artifacts. Therefore, we use the blending strategy [23, 40] and re-design the diffusion process as follows. Instead of converting from noise maps to images in the conventional T2I method, the diffusion process in our DMP approach directly maps between the input image $x$ and output $y$. As illustrated in Figure 2, the proposed diffusion process is formulated as
@@
Parameterization. We explore various parameterizations for the U-Net model to make different predictions. As discussed in Section 4.5, we empirically find that the v-prediction [62] works well for the dense prediction tasks. Specifically, the U-Net model $v_{\theta}$ is fine-tuned using the following mean square loss:
@@ 4.1 Experiment Setup
We evaluate the proposed DMP approach using five dense prediction tasks, including 3D property estimation (i.e., depth and normals), semantic segmentation, and intrinsic image decomposition (i.e., albedo and shading).
@@
Normal Depth
In-domain Out-of-domain In-domain Out-of-domain
@@
SPADE [47] 0.0708 0.1635 0.1268 0.2833 0.2132 0.4961 0.1379 0.3587 0.3190 0.2554
DRIT++ [36] 0.0784 0.1723 0.1350 0.3006 0.3792 0.2458 0.2134 0.4374 0.2585 0.3212
SDEdit [44] 0.2599 0.5087 0.2675 0.5293 0.4656 0.3533 0.3240 0.6640 0.2495 0.3382
DDIB [71] 0.1849 0.4210 0.2271 0.4847 0.3087 0.5130 0.2367 0.6275 0.2788 0.3120
IP2P (hard) [9] 0.3017 0.5468 0.3168 0.5757 0.4834 0.3235 0.3358 0.6450 0.2252 0.3461
IP2P (learned) [9] 0.3550 0.7181 0.3397 0.6836 0.3965 0.3302 0.3494 0.5182 0.2664 0.3261
VISII [46] 0.2081 0.4386 0.2448 0.4895 0.3498 0.4405 0.2912 0.5364 0.2855 0.3181
DMP 0.0514 0.1156 0.0872 0.1886 0.1072 0.8861 0.1020 0.2117 0.6395 0.1360
@@
Datasets. To better analyze generalizability by varying training and test data domains, we conduct the evaluation with synthetic images. We first generate diverse text descriptions using a large language model [45] by filling a keyword in a prompt template modified from the one used in Parmar et al. [50]. We then use a text-to-image diffusion model [57] to synthesize the images according to the text descriptions. Second, we use the following off-the-shelf predictors to generate the pseudo ground truth for each image: Omnidata v2 [29] for surface normals, ZoeDepth [7] for monocular depth, EVA-2 [17] for semantic segmentation, and PIE-Net [14] for intrinsic image decomposition (albedo, shading). Finally, we follow the protocol used in Bhattad et al. [8] to generate a set of training data, and three sets of test data:
@@
  • VISII [46]: We use their approach for fine-tuning with the training set and inference.
@@
Albedo Shading
In Out In Out
SPADE [47] 0.0021 0.0030 0.0031 0.0040
DRIT++ [36] 0.0296 0.0392 0.0309 0.0408
SDEdit [44] 0.0375 0.0471 0.0501 0.0671
DDIB [71] 0.0411 0.0443 0.0403 0.0557
@@
IP2P (learned) [9] 0.0215 0.0250 0.0290 0.0309
VISII [46] 0.0145 0.0246 0.0275 0.0285
DMP 0.0041 0.0064 0.0051 0.0070
Table 4: Effect of different parameterizations. We fine-tune the U-Net model to predict different signals in each reverse diffusion step to get the final outputs: predicting the input image $x$, predicting the output $y$, and v-prediction described in Eq. (3). The experiment is conducted on the surface normal prediction task.
@@
Figure 6: Qualitative comparisons between various parameterizations. We fine-tune the U-Net model to predict different signals in each reverse diffusion step to get the final output: predicting the input image $x$, predicting the output $y$, and v-prediction described in Eq. (3).